# The Hands-Up Problem and How to Deal With It: Secondary School Teachers’ Experiences of Debugging in the Classroom

### v1

@@ Abstract
Debugging is a vital but challenging skill for beginner programmers to learn. It is also a difficult skill to teach. For secondary school teachers, who may lack time or relevant knowledge, honing students’ understanding of debugging can be a daunting task. Despite this, little research has explored their perspectives of debugging. To this end, we investigated secondary teachers’ experiences of debugging in the classroom, with a focus on text-based programming. Through thematic analysis of nine semi-structured interviews, we identified a common reliance on the teacher for debugging support, often embodied by many raised hands. We call this phenomenon the ‘hands-up problem’. While more experienced and confident teachers discussed strategies they use for dealing with this, less confident teachers discussed the generally negative consequences of this problem. We recommend further research into debugging-specific pedagogical content knowledge and professional development to help less confident teachers develop counters to the hands-up problem.
Debugging education,
Secondary teachers,
@@ 1 Introduction
The process of finding and fixing errors in a computer program is one of the most fundamental but challenging skills to learn for the beginner programmer. Learners must have sufficient understanding of several components of the programming language they are using to effectively and reliably resolve errors, which is rarely present in practice. Instead, beginners are commonly observed utilising ineffective debugging strategies (e.g., (Fitzgerald et al., 2008; Liu et al., 2017; Whalley et al., 2023)) which can easily stall progress and evoke emotional distress (Kinnunen and Simon, 2010; Gorson et al., 2022).
This makes debugging difficult to teach, with challenges unique to the secondary school context. Teachers may not be confident in their own programming ability (Yadav et al., 2016a; Royal Society, 2017; Sentance and Csizmadia, 2017) or have to support a large number of students with debugging (Michaeli and Romeike, 2019a). External factors may also reduce the effectiveness of debugging instruction, such as a lack of lesson time (Sentance and Csizmadia, 2017; Michaeli and Romeike, 2019a). This is detrimental to both teachers’ and students’ experiences of teaching and learning debugging in the classroom.
While most debugging education research focuses on the student perspective, such as their debugging behaviours, less research has focused on the teacher. This is particularly true for the K-12 context, where programming is becoming more widely taught (Falkner et al., 2019). In fact, many national curricula stipulate the teaching of text-based programming languages (e.g., (Department for Education, 2013; British Columbia Ministry of Education, 2016; Sysło, 2020; Ripani and Vazquez-Brust, 2023; Skolverket, 2024)), bringing additional debugging-related challenges to teachers. Research must therefore remain close to teachers’ current challenges and needs to be of benefit. This can and should help govern future developments in K-12 debugging education.
This paper reports on an interview study capturing the experiences, perspectives, and observations of secondary computing teachers with respect to debugging in their classrooms, specifically in a text-based programming language. Semi-structured interviews with nine teachers were thematically analysed to explore the role debugging plays in programming lessons. We report on teachers’ current approaches and experiences of teaching students about debugging, their observations of students engaging in the debugging process, and their thoughts on the challenges of teaching it. The research questions for this study are:
@@
The findings from this study can be used to help secondary computing teachers and researchers seeking to improve the teaching and learning of debugging. Our results also give some indication of teachers’ pedagogical content knowledge, which can be supported with further research. We argue that such research into teachers’ experiences, be it debugging or more generally, should be conducted regularly. This will help further developments in computing education research to take place with teachers’ experiences at the forefront.
@@ 2 Related Work
An increasing amount of computing education research is being directed towards the teaching and learning of debugging for beginner programmers. This is for good reason; debugging is a core aspect of programming, difficult to master, and has well-documented challenges. Even with the use of generative AI becoming more prevalent in programming education, debugging is still a vital part of the programming process. Indeed, debugging may become a more emphasised part of the process when using generative AI (Denny et al., 2024). Research into beginner debugging education is therefore important to continue.
2.1 Debugging Education Research
Much debugging education research focuses on the student perspective. Students learning to program naturally have fragile knowledge (Perkins and Martin, 1986) and misconceptions (Sorva, 2021) which often manifest as errors in their programs (e.g., (Du Boulay, 1986; Pea, 1986)). Additionally, students from primary school to university level have been observed employing debugging strategies that do not aid successful debugging, instead embodying a trial and error approach (e.g., (Fitzgerald et al., 2008; Liu et al., 2017; Yan et al., 2021; Whalley et al., 2023)). This makes struggle with debugging likely, which has affective consequences. Students may experience strong negative emotions during periods of difficulty when debugging (Kinnunen and Simon, 2010; Gorson et al., 2022), which dampens their initial experiences of learning to program. Alternatively, students may not be motivated to program at all, meaning any opportunity to debug is ignored.
This research has borne several debugging-specific teaching aids over the last few decades. These include systematic debugging teaching processes (e.g., (Carver and Risinger, 1987; Michaeli and Romeike, 2019b)), debugging-specific tooling (e.g., (Guo, 2013; Luxton-Reilly et al., 2018; Ziemann and Reuß, 2024)), beginner-friendly programming error messages (e.g., (Becker, 2016; Denny et al., 2020; Phung et al., 2023)), and debugging exercises (Kerslake, 2024). The impact of many of these approaches has been summarised in Sun et al. (2024) and Yang et al. (2024), who find these ‘interventions’ have generally positive effects on measures of debugging ability. However, research on many of them is confined to one study, and evidence of wider adoption of what is taught in these interventions is lacking.
@@ 2.2 Teacher Perspectives of Debugging in the Classroom
In comparison, relatively little work has focused on school teachers’ perspectives of debugging, despite them being the main observers of students’ debugging experiences. There are particular difficulties with debugging in school classrooms, where teachers may lack confidence in their knowledge of computing (Yadav et al., 2016a; Royal Society, 2017; Sentance and Csizmadia, 2017), have a lack of time to teach debugging (Sentance and Csizmadia, 2017; Michaeli and Romeike, 2019a), or be inundated with students who need support (Michaeli and Romeike, 2019a). Overall, this lack of research motivates the need for better understanding and improving teachers’ experiences of debugging in the classroom.
Teacher-centric research in debugging education can be split into two areas: teacher-student interactions and studies relating to teachers’ pedagogical content knowledge.
@@ 2.2.1 Teacher-Student Interactions
In a classroom setting, teachers frequently assist students with debugging (Michaeli and Romeike, 2019a). These interactions vary widely, but can generally be viewed through the lens of cognitive apprenticeship (Collins et al., 1991). This is an instructional paradigm for teaching problem-solving processes using the principles of traditional apprenticeship.
@@
Teachers’ use of scaffolding is a major theme in many studies of teacher-student interactions. A popular and general teaching tactic for scaffolding support is to ask students questions, with the content of the questions determining the level of scaffolding provided. Some teachers ask open-ended and generic questions (Nixon et al., 2023; Ryan and DeLiema, 2023; Elliott et al., 2024), providing relatively little scaffolding, while other questions may be far more specific (Yadav et al., 2016b; Flood et al., 2018; Nixon et al., 2023; Elliott et al., 2024; Wachter and Michaeli, 2025). Teachers’ initial questions to students may help the teacher to understand the error(s) in students’ program (Elliott et al., 2024; Wachter and Michaeli, 2025). The more open-ended questions also encourage the articulation of students’ thoughts, another aspect of cognitive apprenticeship important for debugging.
@@
More debugging-specific techniques include modelling the debugging process (Ryan and DeLiema, 2023; Wachter and Michaeli, 2025). Unsurprisingly, students’ debugging strategies can differ from their teacher’s, at which point teachers may decide to tailor their approach to the students’ strategies (Ryan and DeLiema, 2023).
@@
While general techniques used by teachers in their interactions with students are useful to know, the support provided is incredibly contextual. Nixon et al. (2023) frame teacher-student interactions as improvisational (Sawyer, 2004), meaning teachers’ general approaches must be adapted to a specific interaction. Other in-class factors, such as the time remaining in the lesson, number of students in the classroom, and the type of error in question, also influence the support the teacher will provide Elliott et al. (2024). In the same vein, these studies each have their own context, some of which are not representative of a typical classroom environment. Some were conducted in summer schools (Ryan and DeLiema, 2023) or simulations of classrooms (Nixon et al., 2023; Wachter and Michaeli, 2025), so not all findings necessarily reflect teachers’ actual in-class interactions with students.
2.2.2 Pedagogical Content Knowledge
@@
The approaches that teachers adopt when providing debugging support, and the knowledge required to employ them, can be considered part of one’s ‘debugging pedagogical content knowledge’ (PCK) (Shulman, 1986; Carlson et al., 2019; Tsan et al., 2022). Shulman (1986) defines this as ‘the ways of representing and formulating the subject that make it comprehensible to others’ (p. 9). PCK can be developed in several ways: teaching experience, research-informed teaching resources, or professional development (PD) (Shulman, 1986). Indeed, some work has reported the benefits of debugging-specific PD on teachers’ confidence with teaching debugging (Elliott et al., 2022; Michaeli and Romeike, 2022; Nixon et al., 2023).
@@
Despite the volume of pedagogical and content knowledge required to teach debugging effectively, little work has gone into conceptualising a PCK specifically for debugging. One study has analysed school teachers’ level of debugging PCK in Scratch (Tsan et al., 2022), but did not use or develop a formal model. Related work has instead focused on developing computing or programming PCK more generally (Saeli et al., 2011; Hubwieser et al., 2013). However, debugging learning trajectories (Rich et al., 2019; Du et al., 2021) are useful organisations of knowledge that could inform the development of a debugging PCK. Although designed for curriculum development, they could, for example, be used to visualise content knowledge to teachers in an intuitive manner. This would be particularly helpful for non-specialist teachers who need to know the fundamental debugging concepts they need to understand.
@@
Overall, relatively little research has been conducted into teachers’ experiences of debugging in the classroom when compared to student-focused research. Studies on teacher-student debugging interactions indicate that teachers are generally aware of a variety of teaching approaches that can be applied to debugging. However, some of these studies do not take place in the classroom, where teachers’ time for interactions with students may be severely limited (Michaeli and Romeike, 2019a). For less experienced or confident teachers, debugging-specific PD is a helpful way of developing teachers’ debugging PCK, which itself is a concept that could be more formally developed.
@@ 3.1 Participants
Teachers were purposively sampled for the study, with the criteria that they must be teaching a text-based programming language at a lower secondary level [^1] By lower secondary, we refer to the first three years of secondary education, known as key stage three in England, equivalent to grades 6-8 and ages 11-14.. A call for eligible teachers to partake in the study was published in a teacher research network newsletter. Teachers interested in participating were invited to respond, in turn incorporating an element of convenience sampling.
Nine teachers took part in the study, with seven from English and two from non-English countries. Each of these teachers [^2] Apart from one teacher, who was between teaching jobs at the time of interviewing. taught computing as a compulsory subject to their lower secondary students. Teachers from state-funded schools followed the English computing curriculum (Department for Education, 2013) while teachers from private or independent schools were able to implement their own programme of study. Their students would have learnt to program in a block-based programming language in primary school. Each teacher received a bookshop voucher worth £10 as a token of appreciation for participating. Table 1 reports the demographic information of the teachers interviewed.
@@ 3.2 Procedure and Data Collection
We conducted semi-structured interviews with the teachers to allow for ample time and flexibility for teachers to discuss matters relevant to teaching or observations of debugging in their classrooms. Moreover, as per our research questions, we were particularly interested in teachers’ reported feelings and experiences, which suited interviews well.
Interviews took place online from October to December 2023, using either Google Meet or Microsoft Teams, ranging in length from 35-65 minutes. Each teacher was interviewed individually, apart from Chloe and Deborah, who were interviewed together due to time constraints. Before being interviewed, each teacher completed a consent form and demographic questionnaire, in accordance with the study procedure approved by our department’s ethics committee.
The interview questions facilitated two main points of discussion. Firstly, teachers’ own experiences and perspectives with debugging in the classroom. That is, their personal thoughts and recollections related to debugging in programming lessons. This could include teaching approaches they used, general feelings they have experienced about debugging, or beliefs around the role of debugging when teaching programming. Second, we asked teachers about their observations of students’ experiences with debugging, including any common errors, behaviours, or reactions that they commonly see among students. The interview schedule, along with the codebook definitions and agreement procedure, is available to view online [^3] https://github.com/LaurieGale10/secondary-teachers-experiences-debugging-study-materials. We did not have consent from the teachers to publish their interview data.
Each interview was recorded and automatically transcribed. These transcripts were then manually read through in conjunction with the recording to correct any mistakes and remove any mentions of personal or school-specific information, while also being used for initial memoing. These transcripts were sent to teachers within two weeks of the interview, who were then given two weeks to redact or clarify anything they wished to be omitted or corrected. For the teachers who took this opportunity, a final version of the transcript was sent to them for their records. After this period, transcripts were disassociated with the teachers’ contact details.

### v3

@@ Abstract
Debugging is a vital but challenging skill for beginner programmers to learn. It is also a difficult skill to teach. For secondary school teachers, who may lack time or programming experience, honing students’ understanding of debugging can be a daunting task. Despite this, little research has explored their perspectives of debugging. To this end, we investigated secondary teachers’ experiences of debugging in the classroom, with a focus on text-based programming. Through thematic analysis of nine semi-structured interviews, we identified a common reliance on the teacher for debugging support, embodied by many raised hands. We call this phenomenon the ‘hands-up problem’. While more experienced and confident teachers discussed strategies they use to counteract this, less confident teachers discussed the negative consequences of this problem. We recommend further research into debugging-specific pedagogical content knowledge and professional development to help less confident teachers develop approaches for supporting their students with debugging.
@@ 1 Introduction
The process of finding and fixing errors in a computer program is one of the most fundamental but challenging skills to learn for the beginner programmer. Learners must have sufficient understanding of several components of the programming language they are using to effectively and reliably resolve errors, which is rarely present in practice. Instead, beginners are commonly observed utilising ineffective debugging strategies (e.g., (ConditionsLearningNovices; FindingFixingFlailing; ProblemSolvingDebuggingK68)) which can easily stall progress and evoke emotional distress (ProgrammingAssignmentsEmotionalToll; EDAProgrammingEmotions).
Secondary school teachers play a key role in developing their students’ debugging ability. In an ideal classroom environment, they are on hand to teach students effective debugging strategies, support those struggling to debug, and foster positive attitudes towards errors. In practice, this is not always the case, with significant challenges unique to the school context. Teachers may not be confident in their programming ability (ExpandingCSChallengesExperiences; RS2017Report; ComputingInCurriculumChallengesStrategies), lack the lesson time to explicitly teach debugging (ComputingInCurriculumChallengesStrategies; CurrentPerspectivesDebugging), or have to support a large number of students with fixing errors (CurrentPerspectivesDebugging). These challenges are arguably exacerbated when teaching text-based programming languages (TBPLs), which is stipulated in many secondary-level national curricula (e.g., (UKSecondaryComputingCurriculum; BritishColumbiaADSTCurriculum; PolandCSCurriculumChallengesSolutions; ProgrammingSkillsInLatinAmerica; Swedish2024Curriculum)). Although students perceive these languages as more realistic and expressive than block-based languages (ToBlockOrNotToBlock), they have a far larger scope for errors and subsequent struggle with resolving them (JadudCompilationBehaviour; UnderstandingSyntaxBarrier). Teaching with TBPLs also poses an additional layer of difficulty for time-poor or confidence-lacking teachers. These factors can easily make debugging a significant source of negative experiences for students and teachers in secondary school programming lessons.
While most debugging education research has focused on the student perspective, teachers’ classroom experiences are far less studied. This work is always important to conduct; understanding teachers’ experiences is essential for ensuring that research remains close to the challenges teachers are facing in the classroom. Focusing on debugging is particularly important as (text-based) programming becomes more widely taught in schools around the world.
This paper reports on secondary computing teachers’ experiences with the teaching and learning of debugging in a text-based programming language in their classrooms. Through a thematic analysis of semi-structured interviews with nine teachers, we report on teachers’ current approaches for teaching various aspects of debugging, their observations of students engaging in the debugging process, and their challenges with teaching it. The research questions for this study are:
@@
Together with related work, we use our findings to conceptualise a significant in-class challenge for computing teachers that derives from student struggle with debugging, titled the hands-up problem. We introduce this as a representation of a classroom environment unconducive to positive experiences with learning to program in Section 5. We end with recommendations on how the hands-up problem could be alleviated and further developed as a concept. Our findings can be used by school teachers and researchers seeking to understand other teachers’ classroom experiences with debugging and improve the classroom environment of programming lessons.
@@ 2 Related Work
An increasing amount of computing education research is being directed towards the teaching and learning of debugging. This is for good reason: debugging is a core aspect of programming, difficult to master, and has well-documented challenges. Even as generative artificial intelligence becomes more prevalent in programming education, debugging is still, and may be increasingly, a vital aspect of programming (ComputingEducationGenAIEra).
2.1 Student Perspectives of Debugging
Most debugging education research has focused on the student perspective. Students learning to program naturally have fragile knowledge (FragileKnowledgeOriginalPaper) and misconceptions which often manifest as errors in their programs (PeaConceptualBugs; CommonLogicErrors; MisconceptionsBeginnerProgrammer). For students learning a TBPL, the scope for errors is naturally far larger than in BBPLs. Even minor syntactical errors can pose a significant barrier to a working program (JadudCompilationBehaviour; UnderstandingSyntaxBarrier), not helped by unclear programming error messages (ErrorMessageWorkingGroup). These challenges arguably make debugging in TBPLs more difficult than in BBPLs.
When trying to fix these errors, strategies that do not aid successful debugging have been observed among students using block-based (ProblemSolvingDebuggingK68; ElementaryPuzzleBasedDebugging) and text-based (FindingFixingFlailing; ThinkAloudNoviceDebugging) environments. These often embody a cycle of unreflective tinkering (ConditionsLearningNovices), which makes struggle with debugging likely. This struggle can evoke strong negative emotions (ProgrammingAssignmentsEmotionalToll; EDAProgrammingEmotions) and ultimately dampen students’ experiences of learning to program.
The challenges beginner programmers face when debugging have motivated an increased focus on debugging-specific teaching aids over the last few decades. These include systematic debugging processes (e.g., (CarverImprovingChildrensDebugging; SystematicProcessMichaeli)), debugging-specific tooling (e.g., (PythonTutor; Ladebug; SystematicDebuggingLogicalErrors)) and beginner-friendly programming error messages (e.g., (EffectiveApproachErrorMessageEnhancement; ErrorMessageReadability; HighPrecisionFeedbackSyntaxErrors)). Many of these have been specifically designed for TBPLs. Summaries of pedagogical debugging ‘interventions’ have found initial positive effects on measures of debugging ability (DebuggingMetaAnalysis; DebuggingInterventionLitReview), although evidence of wider adoption is lacking.
@@ 2.2 Teacher Perspectives of Debugging in the Classroom
Compared to student-centric research, little work has focused on teachers’ perspectives on teaching debugging. Naturally, students’ experiences with debugging in the classroom are mediated by the teacher, who may employ a range of strategies to develop students’ confidence in resolving errors. However, school teachers face many challenges in providing high-quality debugging instruction. At a subject level, teachers have reported a lack of time and confidence to teach computing (ExpandingCSChallengesExperiences; RS2017Report; ComputingInCurriculumChallengesStrategies). As a consequence, teachers may lack time to explicitly teach debugging or be inundated with students in need of support (CurrentPerspectivesDebugging).
To our knowledge, very few studies have specifically investigated teachers’ experiences of debugging in the classroom, despite the aforementioned challenges. Rather, related work has generally investigated teacher-student interactions or concepts related to pedagogical content knowledge.
@@ 2.2.1 Teacher-Student Interactions
In a classroom setting, teachers frequently assist students with debugging (CurrentPerspectivesDebugging). These interactions vary widely, but can generally be viewed through the lens of cognitive apprenticeship (CognitiveApprenticeshipOriginalArticle), an instructional paradigm for teaching problem-solving processes using the principles of traditional apprenticeship.
@@
Teachers’ use of scaffolding is a major theme in many studies of teacher-student interactions. A popular and general scaffolding technique is to ask students questions, with the content of the questions determining the level of scaffolding provided. Some teachers ask open-ended and generic questions (SupportingStudentsScienceInquiry; ReflectionsSustainedDebuggingSupport; TeacherSupportForDebuggingPhysicalComputing), providing relatively little scaffolding, while other questions may be far more specific (MeasuringCSPCK; InteractionalWorkLearningDebug; SupportingStudentsScienceInquiry; TeacherSupportForDebuggingPhysicalComputing; TeacherDiagnosticInterventionSkills). Teachers’ initial questions to students may help the teacher to understand the error(s) in students’ programs (TeacherSupportForDebuggingPhysicalComputing; TeacherDiagnosticInterventionSkills). More open-ended questions also encourage students to articulate their thoughts, another aspect of cognitive apprenticeship important for debugging.
@@
More debugging-specific techniques include modelling the debugging process (ReflectionsSustainedDebuggingSupport; TeacherDiagnosticInterventionSkills). Unsurprisingly, students’ debugging strategies sometimes differ from their teachers’, at which point teachers may decide to tailor their approach to the students’ strategies (ReflectionsSustainedDebuggingSupport).
@@
While general techniques used by teachers in their interactions with students are useful to know, the support provided is incredibly contextual. SupportingStudentsScienceInquiry frame teacher-student interactions as improvisational, meaning teachers’ general approaches must be adapted to a specific interaction. Other in-class factors, such as the time remaining in the lesson, the number of students in the classroom, and the type of error in question, also influence the support the teacher will provide (TeacherSupportForDebuggingPhysicalComputing). In the same vein, these studies each have their own context, some of which are not representative of an actual classroom environment. Some were conducted in summer schools (ReflectionsSustainedDebuggingSupport) or simulations of classrooms (SupportingStudentsScienceInquiry; TeacherDiagnosticInterventionSkills), so not all findings necessarily reflect teachers’ actual in-class interactions with students.
2.2.2 Pedagogical Content Knowledge and Professional Development
@@
The approaches that teachers adopt when providing debugging support, and the knowledge required to employ them, form part of their ‘debugging pedagogical content knowledge’ (PCK) (PCKOriginalArticle; RefinedConsensusModel). PCKOriginalArticle defines this as “the ways of representing and formulating the subject that make it comprehensible to others” (p. 9). PCK can be developed in several ways: teaching experience, research-informed teaching resources, or professional development (PD) (PCKOriginalArticle). Some debugging-specific PD activities have been developed in previous work, involving frequent teacher reflection (EffectsOfProfessionalDevelopmentK12Debugging) and simulating the student perspective (DoINeedToKnowIfTheTeacher; SupportingStudentsScienceInquiry). Although analyses of these activities only included small cohorts of teachers, they reported benefitting from a low-stakes environment to improve their knowledge of debugging strategies, common errors, and teaching approaches (DoINeedToKnowIfTheTeacher; EffectsOfProfessionalDevelopmentK12Debugging; SupportingStudentsScienceInquiry).
@@
Despite the volume of pedagogical and content knowledge required to teach debugging effectively, we are not aware of a model of PCK specifically for debugging. One study analysed school teachers’ level of debugging PCK (DebuggingPCK), but did not propose a formal model for it and focused on Scratch. Related work has focused on developing computing or programming PCK more generally (ProgrammingPCK; CSPCKConceptualisation). A related concept that could inform the development of a debugging PCK is the idea of debugging learning trajectories (DebuggingLearningTrajectory; DebuggingLearningTrajectoryTextBasedLanguages). Although designed for curriculum development, they could be used to intuitively visualise content knowledge to teachers. This would be particularly helpful for non-specialist teachers wanting to understand fundamental debugging concepts.
@@
Overall, relatively little research has been conducted into teachers’ experiences of debugging in the classroom. Studies on teacher-student debugging interactions indicate that some teachers are aware of a variety of teaching approaches that can be applied to debugging. However, several of these studies do not take place in the classroom, where teachers’ time for interactions with students may be severely limited (CurrentPerspectivesDebugging). For less experienced or confident teachers, debugging-specific PD is a helpful way of developing teachers’ debugging PCK, which itself is a concept that could be further developed.
@@ 3.1 Participants
Teachers were purposively sampled for the study, with the criteria that they must be teaching a text-based programming language at a lower secondary level [^1] By lower secondary, we refer to the first three years of secondary education, known as key stage three in England, equivalent to grades 6-8 and ages 11-14.. Eligible teachers were invited to partake in the study through a teacher research network newsletter, hence incorporating an element of convenience sampling.
Nine teachers took part in the study. Seven of these were from England, where the average class size for state-funded schools is 26 students (UKAverageClassSizes), and two were from other countries. Each of these teachers [^2] Apart from one teacher, who was between teaching jobs at the time of interviewing. taught computing as a compulsory subject to their lower secondary students. Teachers from state-funded and non-English schools followed the English computing curriculum, which mandates the “use two or more programming languages, at least one of which is textual” (UKSecondaryComputingCurriculum, p. 2). Teachers from private or independent schools were able to implement their own programme of study. Python was taught by all teachers, with some also teaching other TBPLs such as Swift. Prior to learning these TBPLs, their students would have been taught a BBPL in primary school (UKPrimaryComputingCurriculum). Each teacher received a bookshop voucher worth £10 as a token of appreciation for participating. Table 1 reports the demographic information of the teachers interviewed.
@@ 3.2 Procedure and Data Collection
We conducted semi-structured interviews to allow teachers ample time and flexibility to discuss matters relevant to their experiences of debugging in the classroom. Moreover, as per our research questions, we were particularly interested in teachers’ reported feelings and experiences, which suited interviews well.
Interviews took place virtually from October to December 2023, using Google Meet or Microsoft Teams. Each teacher was interviewed individually, apart from Chloe and Deborah, who were interviewed together. Before being interviewed, each teacher completed a consent form and demographic questionnaire, in accordance with the study procedure approved by our department’s ethics committee. Interviews lasted between 35-65 minutes.
The interviews facilitated two main points of discussion. Firstly, teachers’ own experiences and perspectives with debugging in the classroom. That is, their personal thoughts and recollections related to debugging in programming lessons. This could include teaching approaches they used, general feelings they have experienced about debugging, or beliefs around the role of debugging when teaching programming. Second, we asked teachers about their observations of students’ experiences with debugging, including any common errors, behaviours, or reactions that they commonly see among students. The interview schedule, along with the codebook definitions and agreement procedure, is available to view in the study repository (StudyRepository). We did not have consent from the teachers to publish the transcripts.
Each interview was recorded and automatically transcribed. The transcripts were then manually read through in conjunction with the recording to correct any mistakes and remove any mentions of personal or school-specific information, while also being used for initial memoing. These transcripts were sent to teachers within two weeks of the interview, who were given two weeks to mention anything they wished to be omitted or corrected. For the teachers who took this opportunity, a final version of the transcript was sent to them for their records. After this period, transcripts were disassociated with the teachers’ contact details.